HUBBLE SPACE TELESCOPE DAILY REPORT #2973 PERIOD COVERED: 0000Z (UTC) 10/09/01 - 0000Z (UTC) 10/10/01 Daily Status Report as of 283/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of STIS/CCD 9285 (POMS Test Proposal: STIS Non-scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.2 Completed Three Sets of STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.3 Completed Three Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.4 Completed FGS//1 9308 (Verifying FGS FSW on Faint Targets) Fine Guidance Sensor #1 was used to verify new FGS flight software that has been implemented to facilitate the reliable acquisition and tracking of faint stars in POS mode and for the reliable positioning of the scan path in Trans mode. There were no problems. 1.5 Completed STIS/CCD/MA1/MA2 8487 (High Resolution Observations of HD 195965) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to observe HD 195965 at high resolution to determine the velocity structure and abundances of the interstellar gas along the sight line. The object has been observed with FUSE, and the investigators seek understand the detailed nature of the ISM to gain insight into the D/H ratio along the sight line. No problems were noted. 1.6 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. There were no problems. 1.7 Completed WF/PC-2 9042 (An Archive To Detect The Progenitors Of Massive, Core-Collapse Supernovae) The WF/PC-2 was used to search for supernovae which have massive star progenitors. The already extensive HST archive and high-resolution ground-based images of galaxies within ~20 Mpc enables us to resolve and quantify their individual bright stellar content. As massive, evolved stars are the most luminous single objects in a galaxy, the progenitors of core-collapse supernovae should be directly detectable on pre-explosion images. One Type II progenitor has been observed this year, and the investigators have proposed a short, companion WFPC2 proposal to confirm this candidate and identify a second. The observations completed nominally. 1.8 Completed Two Sets of STIS/MA2 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.9 Completed STIS/CCD 9066 (Closing in on the Hydrogen Reionization Edge of the Universe) The Space Telescope Imaging Spectrograph (CCD) was used in parallel constrain the Hydrogen reionization edge in emission that marks the transition from a neutral to a fully ionized IGM at a predicted redshifts. As described in 2.1 and HSTAR 8363, the acquisition for this proposal defaulted to fine lock backup on one FGS only, possibly affecting the proposal. Otherwise, the proposal completed uneventfully. 1.10 Completed WF/PC-2 8882 (Chandra and HST Observations of the Brightest Cluster Lenses) The WF/PC-2 was used in conjunction with Chandra to perform observations of the four X-ray brightest, cooling-flow clusters with bright gravitational arcs, newly identified from the ROSAT XBACS/BCS lensing study. As described in 2.1 and HSTAR 8363, the acquisition for this proposal defaulted to fine lock backup on one FGS only, possibly affecting the proposal. Otherwise, the observations completed with no other reported problems. 1.11 Completed WF/PC-2 8815 (Cycle 9 Earth Flats) The WF/PC-2 was used to monitor flatfield stability. This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. As described in 2.1 and HSTAR 8363, the acquisition for this proposal defaulted to fine lock backup on one FGS only, possibly affecting the proposal. Otherwise, the proposal completed nominally. 1.12 Completed WF/PC-2 9057 (Host Galaxies of Obscured QSOs Identified by 2MASS) The WF/PC-2 was used to perform a snapshot survey of red QSOs discovered in The Two Micron All Sky Survey {2MASS} to investigate the detailed properties of their host galaxies. This large, possibly dominant, population of QSOs in the local universe has been previously overlooked because reddening by {intrinsic} obscuration along our line of sight causes their colors to be too red for identification by traditional "UV- excess" techniques. Their near-IR colors are similar to PG- type {UV-excess} QSOs, but it is far from certain whether they are indeed from the same parent population or represent a completely new class of QSO. There were no reported problems. 1.13 Completed STIS/CCD 9186 (D/H in Lyman Limit Absorbers with Simple Velocity Structure) The Space Telescope Imaging Spectrograph (CCD) was used to observe the deuterium abundance in QSO absorbers that provides a direct measurement of the baryonic density of the Universe. This proposal will obtain new high resolution data of two carefully selected, intermediate redshift absorption systems. No problems were encountered. 1.14 Completed WF/PC-2 8936 (Cycle 10 Supplemental Darks Pt1/3) The WF/PC-2 was used to perform a dark calibration program that obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 5 Per HSTAR 8363, the acquisition at 282/095014Z defaulted to fine lock backup on FGS-2 only when the scan step limit was exceeded on FGS-3. The proposals detailed in 1.9, 1.10, and 1.11 may have been affected. Scheduled Re-acquisitions: 10 Successful: 10 2.2 FHST Updates: Scheduled: 11 Successful: 11 2.3 Operations Notes: The SSR EDAC error counter was cleared three times per ROP SR-1A. A TTR was written when there was a required re-transmit at 282/2029Z during a NSSC-1 load. Accordingly, the SI C&DH errors were reset at 282/2033Z using ROP NS-5. The STIS MCE-2 reset at 283/020944Z while the low voltage was on and while inside an SAA interval. The SIS flight software error counter was reset at 283/0219Z as directed by ROP NS-12. MAMA-2 was recovered at 283/0441Z, the time of the next high voltage on commanding, via normal SMS commanding. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW